### v1

@@ 4 Experiments
Our evaluations focus on multi-turn conversational workloads. We use LMsys-Chat-1M zheng2024lmsyschat1mlargescalerealworldllm and ShareGPT zheng2023judging as our datasets, which consist of large-scale, real-world human-assistant conversations collected from ChatGPT and LMsys platforms.
@@
We collect several key metrics for the prefill stage, including TTFT, P90 latency, average request per second (RPS), and SLO violation rate wang2024revisiting.
Figure 6: Comparison of PLA-Serve and SGLang on one H200 node. The top two lines of figures correspond to the temporal disaggregation setting (with prefill instance of 1), while the bottom two lines of figures correspond to the spatial disaggregation setting (with prefill instance of 8). We report RPS, average latency, and P90 latency under four configurations: Vanilla SGLang PD disaggregation (blue line), PLA (only CUDA Graphs enabled, orange line), PLA (only disaggregation enabled, green line), and Full PLA (red line).
@@
We compare PLA-Serve against SGLang and vLLM, both are state-of-the-art LLM serving systems. SGLang is a widely adopted serving system in both academia and industry; it implements continuous batching to improve throughput and radix-attention zheng2024sglangefficientexecutionstructured to mitigate memory fragmentation during KV-cache allocation. However, neither SGLang nor vLLM supports CUDA Graph during the prefill phase, and their batching policies rely solely on available memory capacity, so they cannot adjust their batching strategies according to the workload characteristics.
@@ 4.1 Numerical Results and Analysis
Figure 6 compares the performance of PLA-Serve with SGLang (with PD disaggregation) and its two partial variants (PLA with CUDA Graph only and PLA with Disaggregation only) under a sustained client load with varying concurrency levels (from 1 to 64). The requests are drawn from real multi-turn conversations in the ShareGPT-4 dataset. We evaluate three models, Qwen2.5-7/14/32B, under both the single-instance (temporal disaggregation) and 8-instance (spatial disaggregation) settings.
PLA consistently outperforms SGLang and its two partial variants across all key metrics: RPS, average latency, and P90 latency. The benefits of PLA’s scheduling mechanism and CUDA Graph optimization become more pronounced under high concurrency. Specifically, PLA achieves up to 20% and 33% higher RPS than the baseline in the single-prefill instance and 8-prefill instance settings, while reducing average latency by 20% respectively.
@@
Figure 7: SLO violation rate under varying client concurrency levels using the LMsys-Chat-1M dataset. Results are shown for PLA-Serve, SGLang (PD disaggregation), SGLang (PD disaggregation with router), and vLLM (PD disaggregation) under two settings: (top) single-instance (temporal disaggregation) and (bottom) 8-instance (spatial disaggregation).
In Figure 7, we use the LMsys-Chat-1M dataset and assume that request arrivals follow a Poisson process with an average arrival rate $\lambda$, while each request’s service time follows an empirical distribution measured from model execution. We set the TTFT SLO to 0.4s and vary the client-side concurrency levels to observe the actual SLO violation rate. SGLang supports data parallel (DP) serving based on a router that dispatches requests to different workers using either round-robin or load-balancing strategies; however, the router is unaware of the SLOs of individual requests. As shown in the figure, within a single instance, PLA-Serve reduces the SLO violation rate by approximately 10% compared with SGLang (PD disaggregation) with router, and by around 30% compared with Vanilla DP. Under the 8-instance spatial disaggregation setting, PLA-Serve achieves zero SLO violations, whereas SGLang with router still exhibits a 4.7% violation rate.
In Figure 8, we evaluate the compatibility of PLA under non-PD-disaggregated settings by mixing prefill and decode requests at different concurrency levels. In both single-instance and multi-instance configurations, the request-per-second (RPS) of prefill requests decreases under the Mix with Decode condition, indicating that PLA can fully exploit the throughput benefits of CUDA Graphs for short prefill requests only within the PD-disaggregated architecture. When mixed with decode workloads, the lack of continuous batching introduces additional inter-batch latency, resulting in degraded overall performance.
As shown in Table 2, we evaluate the distillation task where each request has no strict deadline (i.e., the waiting window can be relatively large). Under this setting, with four prefill instances and a decoding length of 1K tokens, PLA achieves about an 8% reduction in time consumption compared to SGLang (Vanilla PD disaggregation).
@@ 4.2 Cost analysis
In PLA deployment, CUDA Graphs are captured into memory during initialization. Each graph is bound to a fixed kernel configuration and cannot adapt to dynamic kernel sizes, so multiple graphs must be captured to cover different token lengths and batch sizes. Each prefill step introduces lookup and selection overhead, and thus, the number of graphs must be limited to balance memory usage and performance. We measure single-graph sizes of 228 MB, 240 MB, and 277 MB for the 7B, 14B, and 32B models, showing that graph size is largely insensitive to model scale. When the system is initialized for the first time, it needs to capture kernels and the KV-cache operations layer by layer, which introduces a certain startup overhead. Experiments show that capturing a single prefill graph incurs an initialization overhead of approximately 8-12 seconds.
@@ 5 Conclusion
In this paper, we propose PLA-Serve, a prefill-length-aware LLM serving system built on the PD disaggregation paradigm to optimize heterogeneous multi-turn conversational workloads. By separating long- and short-prefill requests, PLA-Serve eliminates compute–memory interference in the prefill stage. Its adaptive scheduler (AWD) and CUDA Graph–based execution improve batching efficiency and reduce short-prefill latency. Supporting both temporal and spatial disaggregation, PLA-Serve scales across single- and multi-prefill-instance deployments. Experiments on real-world datasets show that PLA-Serve achieves higher throughput and lower latency than state-of-the-art frameworks (e.g., SGLang and vLLM under PD disaggregation), demonstrating its effectiveness under high concurrency.

### v2

@@ 4 Experiments
Our evaluations focus on multi-turn conversational workloads. We use LMsys-Chat-1M Zheng et al. (2024a) and ShareGPT Zheng et al. (2023) as our datasets, which consist of large-scale, real-world human-assistant conversations collected from ChatGPT and LMsys platforms.
@@
We collect several key metrics for the prefill stage, including TTFT, P90 latency, average request per second (RPS), and SLO violation rate Wang et al. (2024).
Figure 6: Comparison of LAPS and SGLang on one H200 node. The top two lines of figures correspond to the temporal disaggregation setting (with prefill instance of 1), while the bottom two lines of figures correspond to the spatial disaggregation setting (with prefill instance of 8). We report RPS, average latency, and P90 latency under four configurations: Vanilla SGLang PD disaggregation (blue line), LAPS (only CUDA Graphs enabled, orange line), LAPS (only disaggregation enabled, green line), and Full LAPS (red line).
@@
We compare LAPS against SGLang and vLLM, both are state-of-the-art LLM serving systems. SGLang is a widely adopted serving system in both academia and industry; it implements continuous batching to improve throughput and radix-attention Zheng et al. (2024b) to mitigate memory fragmentation during KV-cache allocation. However, neither SGLang nor vLLM supports CUDA Graph during the prefill phase, and their batching policies rely solely on available memory capacity, so they cannot adjust their batching strategies according to the workload characteristics.
@@ 4.1 Numerical Results and Analysis
Figure 6 compares the performance of LAPS with SGLang (with PD disaggregation) and its two partial variants (LAPS with CUDA Graph only and LAPS with Disaggregation only) under a sustained client load with varying concurrency levels (from 1 to 64). The requests are drawn from real multi-turn conversations in the ShareGPT-4 dataset. We evaluate three models, Qwen2.5-7/14/32B, under both the single-instance (temporal disaggregation) and 8-instance (spatial disaggregation) settings.
LAPS consistently outperforms SGLang and its two partial variants across all key metrics: RPS, average latency, and P90 latency. The benefits of LAPS’s scheduling mechanism and CUDA Graph optimization become more pronounced under high concurrency. Specifically, LAPS achieves up to 20% and 33% higher RPS than the baseline in the single-prefill instance and 8-prefill instance settings, while reducing average latency by 20% respectively.
@@
Figure 7: SLO violation rate under varying client concurrency levels using the LMsys-Chat-1M dataset. Results are shown for LAPS, SGLang (PD disaggregation), SGLang (PD disaggregation with router), and vLLM (PD disaggregation) under two settings: (top) single-instance (temporal disaggregation) and (bottom) 8-instance (spatial disaggregation).
In Figure 7, we use the LMsys-Chat-1M dataset and assume that request arrivals follow a Poisson process with an average arrival rate $\lambda$, while each request’s service time follows an empirical distribution measured from model execution. We set the TTFT SLO to 0.4s and vary the client-side concurrency levels to observe the actual SLO violation rate. SGLang supports data parallel (DP) serving based on a router that dispatches requests to different workers using either round-robin or load-balancing strategies; however, the router is unaware of the SLOs of individual requests. As shown in the figure, within a single instance, LAPS reduces the SLO violation rate by approximately 10% compared with SGLang (PD disaggregation) with router, and by around 30% compared with Vanilla DP. Under the 8-instance spatial disaggregation setting, LAPS achieves zero SLO violations, whereas SGLang with router still exhibits a 4.7% violation rate.
In Figure 8, we evaluate the compatibility of LAPS under non-PD-disaggregated settings by mixing prefill and decode requests at different concurrency levels. In both single-instance and multi-instance configurations, the request-per-second (RPS) of prefill requests decreases under the Mix with Decode condition, indicating that LAPS can fully exploit the throughput benefits of CUDA Graphs for short prefill requests only within the PD-disaggregated architecture. When mixed with decode workloads, the lack of continuous batching introduces additional inter-batch latency, resulting in degraded overall performance.
As shown in Table 2, we evaluate the distillation task where each request has no strict deadline (i.e., the waiting window can be relatively large). Under this setting, with four prefill instances and a decoding length of 1K tokens, LAPS achieves about an 8% reduction in time consumption compared to SGLang (Vanilla PD disaggregation).
@@ 4.2 Cost analysis
In LAPS deployment, CUDA Graphs are captured into memory during initialization. Each graph is bound to a fixed kernel configuration and cannot adapt to dynamic kernel sizes, so multiple graphs must be captured to cover different token lengths and batch sizes. Each prefill step introduces lookup and selection overhead, and thus, the number of graphs must be limited to balance memory usage and performance. We measure single-graph sizes of 228 MB, 240 MB, and 277 MB for the 7B, 14B, and 32B models, showing that graph size is largely insensitive to model scale. When the system is initialized for the first time, it needs to capture kernels and the KV-cache operations layer by layer, which introduces a certain startup overhead. Experiments show that capturing a single prefill graph incurs an initialization overhead of approximately 8-12 seconds.
@@ 5 Conclusion
In this paper, we propose LAPS, a prefill-length-aware LLM serving system built on the PD disaggregation paradigm to optimize heterogeneous multi-turn conversational workloads. By separating long- and short-prefill requests, LAPS eliminates compute–memory interference in the prefill stage. Its adaptive scheduler (AWD) and CUDA Graph–based execution improve batching efficiency and reduce short-prefill latency. Supporting both temporal and spatial disaggregation, LAPS scales across single- and multi-prefill-instance deployments. Experiments on real-world datasets show that LAPS achieves higher throughput and lower latency than state-of-the-art frameworks (e.g., SGLang and vLLM under PD disaggregation), demonstrating its effectiveness under high concurrency.